find the inverse laplace transform of F (S)=10 (S+1)/S (S+2)(S+3)^3


1
Expert's answer
2021-08-16T14:45:28-0400
"\dfrac{10(s+1)}{s(s+2)(s+3)^3}=\dfrac{A}{s}+\dfrac{B}{s+2}"

"+\dfrac{C}{s+3}+\dfrac{D}{(s+3)^2}+\dfrac{E}{(s+3)^3}"

"10(s+1)=A(s+2)(s+3)^3+Bs(s+3)^3"

"+Cs(s+2)(s+3)^2+Ds(s+2)(s+3)"

"+Es(s+2)"

"s=0,10=A(2)(3)^3=>A=\dfrac{5}{27}"

"s=-2,-10=B(-2)(1)^3=>B=5"

"s=-3,-20=E(-3)(-1)=>E=-\dfrac{20}{3}"

"s^4, 0=A+B+C=>C=-\dfrac{140}{27}"

"s^3, 0=A(9+2)+9B+C(6+2)+D"

"=>D=-\dfrac{50}{9}"

"\dfrac{10(s+1)}{s(s+2)(s+3)^3}=\dfrac{5}{27}\cdot\dfrac{1}{s}+5\cdot\dfrac{1}{s+2}"

"-\dfrac{140}{27}\cdot\dfrac{1}{s+3}-\dfrac{50}{9}\cdot\dfrac{1}{(s+3)^2}-\dfrac{20}{3}\cdot\dfrac{1}{(s+3)^3}"


"L^{-1}(\dfrac{10(s+1)}{s(s+2)(s+3)^3})=\dfrac{5}{27}H(t)+5e^{-2t}"

"-\dfrac{140}{27}e^{-3t}-\dfrac{50}{9}te^{-3t}-\dfrac{10}{3}t^2e^{-3t}"